OpenStudy (anonymous):

A certain kind of passenger jet can travel at a speed of 8.4 × 10^5 meters in 1 hour. At this speed, how far can the jet travel in 5 hours? A. 4.2 × 10^4 meters B. 4.2 × 10^5 meters C. 4.2 × 10^6 meters D. 8.4 × 10^10 meters <----- @shamim

OpenStudy (shamim):

8.4*10^5*5=?

OpenStudy (anonymous):

4200000 Also sorry I typed that wrong. :) I think it's B. :) @shamim

OpenStudy (anonymous):

Also can you check 7 more? Please!!! o_o

OpenStudy (anonymous):

@shamim Am I right?

OpenStudy (shamim):

c

OpenStudy (shamim):

8.4*5=?

OpenStudy (anonymous):

42

OpenStudy (shamim):

u r right

OpenStudy (anonymous):

So which is it?

OpenStudy (shamim):

so ur answer is 42*10^5 =4.2*10^6

OpenStudy (shamim):

is not it C

OpenStudy (anonymous):

OK. :)

OpenStudy (shamim):

anyway i always enjoy to solve ur math